OpenStudy (anonymous):

FAN AND MEDAL: Lucy's goal for her cycling class at the gym is to burn 450 calories in one hour. The number of calories (c) she actually burns in one hour varies no more than 45 calories. Which inequality below represents this scenario? |c - 450| less than or greater to 45 |c + 450| greater than or equal to 45 |c - 45| less than or greater to 450 |c - 45| greater than or equal to 450

OpenStudy (anonymous):

@undeadknight26 @jhonyy9 @wolf1728 @Frostbite @Whitemonsterbunny17 @Taylor<3sRin @HelpBlahBlahBlah Can any of y'all help me? :D

OpenStudy (anonymous):

@OrangeMaster @sammixboo

OpenStudy (anonymous):

Someone help this Girl @Hero and @undeadknight26

OpenStudy (anonymous):

@ranga

undeadknight26 (undeadknight26):

YOu lost me at girl XD I will see what i can do :D

OpenStudy (anonymous):

Thank you!

undeadknight26 (undeadknight26):

I would say A :D

OpenStudy (anonymous):

Can you please explain how you got that answer? @undeadknight26

OpenStudy (anonymous):

oh ur welcome @kaylamarie101

undeadknight26 (undeadknight26):

lolz i googled it XD

OpenStudy (anonymous):

LOL undead

undeadknight26 (undeadknight26):

It is actually quite self explanatory...

OpenStudy (anonymous):

lol wow. Okay, well thanks anyway!

undeadknight26 (undeadknight26):

You put this "Lucy's goal for her cycling class at the gym is to burn 450 calories in one hour. The number of calories (c) she actually burns in one hour varies no more than 45 calories." Into a formula to find C...

undeadknight26 (undeadknight26):

Hope i helped and learn on!

OpenStudy (anonymous):

You did help. Actually do you think that you could check this one out: After a discount of 20%, the price of fitness equipment, including tax, was $73.53. The following equation can be used to find x, the price of the equipment without the discount: x - 0.2x + 0.07(x - 0.2x) = 73.53 What is the price of the fitness equipment without the discount? $74.39 $79.06 $85.90 $95.49

OpenStudy (anonymous):

@undeadknight26

undeadknight26 (undeadknight26):

(x - 0.2x) (1 + 0.07) = 73.53 0.8x (1.07) = 73.53 0.856x = 73.53 x = 73.53 / 0.856 = 85.90 I think...your answer should be 85.90...There an answer :p

OpenStudy (anonymous):

OMG Thank You SO much! Your awesome! :D

OpenStudy (anonymous):

@undeadknight26

undeadknight26 (undeadknight26):

Omg thank you XD Are you in 9th by any chance? FLVS or OCVS?

OpenStudy (anonymous):

lol I'm taking Algebra 1 on FLVS...how'd you know?

undeadknight26 (undeadknight26):

Because i just finished that course lolz...I am in OCVS 9th...just finished it not too long ago XD just lazy today to answer questions due to gaming XD

OpenStudy (anonymous):

haha Wow. What a coincidence (: I am in 11th grade, going into 12th this August but i got a D in the course my 9th grade year, so I am retaking it for grade and GPA recovery lol

OpenStudy (anonymous):

i fell off my chair then while getting up i hit my head talk about bad carma Lolz

undeadknight26 (undeadknight26):

Lol wow XD If you need any help just ask me...i think the assignments might be different though.

OpenStudy (anonymous):

Lol Wow @Hotchellerae21 How'd that happen? And thanks @undeadknight26

OpenStudy (anonymous):

lol Wow